OpenStudy (anonymous):

The intensity of light hitting an object, L, varies inversely with the square of the distance, d, between the light source and the illuminated object. Which of the following expresses this mathematically? (The term k is a constant.)

OpenStudy (anonymous):

A. Ld^2=k B. L=d/k or C. L=kd^2

OpenStudy (rock_mit182):